Fermi National Laboratory
 

Accelerator Update for February 20, 2006 - February 22, 2006

 

Monday February 20

The day shift began with Operations monitoring the Recycler stash, with D-Zero (D0) in access, and with NuMI, MiniBooNE, and the Meson MTest and MCenter experiments taking beam.

At 8:30 AM, a Booster power supply (DOGL3) tripped off. An expert investigated.

Operations turned off beam to MTest at 9:43 AM. The experiment will be in access all day.

D0 came out of access at 9:57 AM.

Operations put the Tevatron (TeV) through a dry squeeze at 10:17 AM.

Operations began preparing the Recycler to stash antiprotons at 12:40 PM. The stash ended at 1:15 PM.

At 1:15 PM, Operations reported that the TeV suffered a sector A1 quench.

Operations prepared Pbar for an access at 1:16 PM.

Operations put the TeV through a dry squeeze at 3 PM.

Pbar came out of access and resumed stacking at 3:03 PM.

At 3:18 PM, NuMI tripped off for the first of two times in ten minutes. It wouldn’t reset after the second trip. Operators investigated and later reported that beam position problems in the Main Injector (MI) had caused the NuMI trips. Experts began working on this problem at the end of the day shift.

At 3:45 PM, a Debuncher power supply (D:IB) began tripping off. An expert investigated and began repairs.

Operations put TeV into shot setup at 3:48 PM.

NuMI resumed taking beam at 5:09 PM.

At 5:32 PM, just as a power supply expert was finishing his repairs on D:IB, a Pbar expert reported that they had to access the tunnel again. Operators prepared the area for access.

Pbar came out of access at 6:14 PM. Operation had a problem turning on an Accumulator power supply (A:IB). Operators investigated and reported a contactor problem. Operations called back in the power supply expert.

Operations established store 4658 at 6:19 PM with an initial luminosity of 142.59E30. MCenter resumed taking beam.

All Pbar work was completed. Pbar resumed stacking at 9:04 PM.

At 10:39 PM, Operations reported that Pbar was stacking at a very poor rate. An operator contacted an expert for help. With expert advice, operators repaired a RF station (DRF1). Normal stacking resumed around 11:30 PM.

Tuesday February 21

The midnight shift began with Operations monitoring store 4658 and the stash, with Pbar stacking, and with NuMI, MiniBooNE, and MCenter taking beam.

The Recycler began preparing to stash at 3:56 AM. The stash ended. Pbar resumed stacking at 4:56 AM. MCenter resumed taking beam.

At 5:30 AM, Operations reported that store 4658 aborted due to a problem with the TeV sector A2 vacuum. Operators investigated and, with the help of an expert, discovered a bad CIA crate power supply.

Operations turned off NuMI at 8 AM for a scheduled access.

The Recycler began preparing to stash at 8:28 AM. The stash ended and MCenter resumed taking beam at 9:52 AM.

The TeV studies ended and Operations prepared D-Zero (D0) for a collision hall access at 2:31 PM.

At 3:03 PM, the MiniBooNE Horn began tripping off. Operations informed experts.

The Recycler began preparing to stash at 3:17 PM. The stash ended around 4 PM.

Operations put the TeV into shot setup at 5:28 PM.

NuMI resumed taking beam at 5:49 PM.

Operations established store 4664 at 8:20 PM with an initial luminosity of 52.44E30. Four antiproton transfers were lost during shot setup due to a timeline problem. MCenter resumed taking beam.

MiniBooNE experts reported at 8:40 PM that the Horn trips were due to a waveform circuit board. A spare board resolved the problem.

The Recycler began preparing to stash at 11:06 PM.

Wednesday February 22

The midnight shift began with Operations monitoring store 4664 and the stash, with Pbar stacking, and with NuMI, MiniBooNE, MTest, and MCenter taking beam.

The stash ended. Pbar, MTest, and MCenter resumed taking beam.

The Recycler began preparing to stash at 3:03 AM. The stash ended. Pbar, MTest, and MCenter resumed taking beam around 3:55 AM.

The Recycler began preparing to stash at 7:14 AM. The stash ended. Pbar, MTest, and MCenter resumed taking beam around 7:55 AM.

The Plans for Wednesday

The plans for today are to end the store and allow D0 and CDF to access their collision halls around 10 AM. After the access, the Recycler will stash. The next TeV shot will occur around 12:30 PM. At 8 AM on Thursday, 2-23-06, Operations will lower MI beam intensity in preparation for the shut down.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/22/2006    email Fermilab